OpenStudy (anonymous):

let f(x) = x^2. Find x such that f(x)=15

OpenStudy (lgbasallote):

f(x) = x^2 and f(x) = 15 therefore x^2 = 15 now solve for x

OpenStudy (lgbasallote):

dont forget square root equals \(\pm\)